Crew 195 Daily Summary Report 23.05.2018
MDRS Daily Summary Report for sol 3
Sunny and Hot- cooking on Mars.
Sol Activity Summary:
Second EVA to Gateway of Lith. Successful sim of medical resuscitation
Look Ahead Plan: EVA to west ridge tomorrow.
Anomalies in work:
ATV 300- continued issues with battery. Unable to hold charge.
Weather:Warmer today. Simulated heat illness
Crew Physical Status:Doing well.
Reports to be file:
Photo of the day